in 180 minutes, the activity of a certain radioactive substance falls to 1/8 its original value. calculate its half life


1
Expert's answer
2021-06-15T14:57:05-0400

"\frac18N_0=N_02^{-\frac{t}{T_{1/2}}}," "\implies"

"T_{1/2}=\frac t 3=\frac{180}{3}=60~min."